if 7% is taken away from $330 how much is left?
Answer by jim_thompson5910(35256)

7% of 330 = 0.07*330 = 23.10

7% of $330 = $23.10

-------------------------------------------------------

So 23.10 is taken away: 330 - 23.10 = 306.90

Leaving $306.90